# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

### v2

@@ III-B3 Rejection Sampling
To ensure the quality and diversity of the negative samples, we apply a series of filters to the candidates. First, validity filtering is performed, which includes deduplication, removal of null values, and elimination of samples that are identical to or contain the ground truth. Subsequently, similarity filtering based on BLEU scores is used to remove negative samples that are too similar to the ground truth.
@@
After the screening process, we retain up to 3 negative samples for each original data point, which constitute the paired dataset for preference alignment training. At the same time, to suppress the model’s tendency to simply repeat the context, we randomly sampled 10% of the data from rejection sampling where the expected completion does not start with the first line of the suffix. For this subset, we constructed suffix-repetition suppression pairs where the positive example is the ground-truth completion and the negative example is the first line of the suffix code. This data was then added to the synthesized dataset. Similarly, we also added 1% of prefix-repetition suppression data.
@@ III-C Model Post-training
Using the data synthesized with the methods described above, we conducted two stages of post-training on the Seed-Coder-8B-Base model [9]: fine-tuning and alignment. Seed-Coder-8B-Base is an 8B-parameter open-source large language model for code developed by the ByteDance Seed team. It serves as the base model for the Seed-Coder series and is designed to provide a powerful starting point for various downstream tasks, including code generation and reasoning. Thus, we use Seed-Coder-8B-Base as our based model for code completetion task.
@@ III-C1 Finetune Stage
The data preprocessing for curriculum learning fine-tuning is primarily divided into three steps:
  • Step 1: Code Parsing.
    After initial filtering, we use Tree-sitter to parse code files and generate Abstract Syntax Trees (ASTs).
  • Step 2: Complexity Sorting.
    We measure complexity by the number of identifier nodes in the AST and sort them in descending order. Symbols are extracted by recursively traversing the AST.
  • Step 3: Sample Selection.
@@
In the curriculum learning fine-tuning stage, we trained the model for one epoch using the AdamW optimizer, with a learning rate set to 3e-6 combined with a 50-step warmup. The training data selected was the top 30% (k=0.3) of the original dataset sorted by complexity, and the input sequences were uniformly truncated to within 8000 tokens. The entire process occupied 4 H800 80G PCIE GPUs, with a global batch size of 64 (micro-batch size of 1). Strategies such as DeepSpeed ZeRO-1, Flash Attention, and gradient checkpointing activation re-computation were used to improve training efficiency. Finally, the process accounted for 800 steps and finished in 20 hours.
@@ III-C2 Alignment Stage
In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64. The model was trained for one epoch with a maximum learning rate of 5e-5, a linear learning rate warmup of 48 steps, and a cosine learning rate decay scheduler. The maximum length for training data was 8192. For LoRA, the rank and alpha were set to 32 and 64, respectively, and the LoRA training modules included various modules within Self-Attention and the feed-forward networks. The training was conducted on 4 NVIDIA H800 80G PCIE GPUs using DeepSpeed ZeRO-2 and a CPU offloading strategy. The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 2. The alignment process lasted for 400 steps within 24 hours, and it can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable.
@@ IV Evaluation
We conducted a comprehensive and rigorous evaluation of SynthCoder and baseline models on both file-level and repository-level code completion tasks. Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities that were released between 2024 and the first half of 2025. Models released earlier, such as CodeLlama-7b-hf [^11] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded. Additionally, we excluded models not specifically fine-tuned for FIM tasks, such as MagicCoder [^12] https://huggingface.co/ise-uiuc/Magicoder-S-DS-6.7B. All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1131313https://github.com/vllm-project/vllm. The final set of models used for comparison is as follows.
  • aiXcoder-7b-v2141414https://huggingface.co/aiXcoder/aixcoder-7b-v2
    A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
  • StarCoder2-7B [^15] https://huggingface.co/bigcode/starcoder2-7b
    A 7B parameter code LLM developed by the BigCode project. It was trained on The-Stack-V2, a large-scale dataset containing multiple programming languages.
  • DeepSeek-Coder-6.7B (-base [^16] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^17] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct)
    A pair of 6.7B code LLMs developed by DeepSeek. The former is the base model, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
  • Qwen2.5-Coder-7B [^18] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^19] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct)
    A pair of 7B code LLMs developed by Alibaba, designed for code understanding and generation. Qwen2.5-Coder-7B is the base model, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version of Qwen2.5-Coder-7B.
  • Seed-Coder-8B (-Base [^20] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^21] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct)
    A pair of 8B parameter code LLMs developed by the ByteDance. The model was trained using a model-driven data selection method, aiming to efficiently enhance its code generation abilities, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
@@
    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^24] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base. Both of them utilize a Mixture-of-Experts (MoE) architecture, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version of DeepSeek-Coder-V2-Lite-Base.
@@
    An efficient instruction-tuned code model from Alibaba’s Qwen3 series, which also employs a Mixture-of-Experts (MoE) architecture and possesses excellent instruction-following capabilities.
To ensure the rigor of our experimental results, we only included models in our comparison that were evaluated in the aforementioned file-level FIM code completion benchmarks and for which the official developers provided a specific FIM prompt format in their technical reports or open-source repository documentation.
  • The Qwen-Coder series (including Qwen2.5-Coder and Qwen3-Coder) models follow the format specified in their technical report [8];
  • The DeepSeek-Coder series models follow the example provided in the official repository [^26] https://github.com/deepseek-ai/DeepSeek-Coder;
  • The Seed-Coder series models follow the official response in an issue on their official open-source repository [^27] https://github.com/ByteDance-Seed/Seed-Coder/issues/4;
  • The StarCoder2 model follows the format described in its technical report [35];
  • The aiXcoder-7B-v2 model’s technical report [10] and open-source repository [^28] https://github.com/aixcoder-plugin/aixcoder-7b-v2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context. Therefore, we only evaluated this model in the file-level code completion tasks.
To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS. To validate the impact of different training stages and data compositions, we evaluated three versions of our model: SynthCoder-8B-SFT, SynthCoder-8B-DPO-wo-RS, and SynthCoder-8B-DPO-RS (the final version, referred to as SynthCoder-8B). We compared these against the base model, Seed-Coder-8B-Base. The differences between these model versions are detailed in Table I. To save space, all numerical values in tables are rounded to one decimal place. The percentage of increase or decrease refers to the change relative to SynthCoder-8B-DPO-RS.
@@
SynthCoder-8B-DPO-RS YES YES YES
IV-A File Level Completion
We select aiXcoder dataset as the evaluation dataset to verify our model’s capacity on file level code completetion.
@@ IV-A File Level Completion
We utilized the dataset [^29] https://github.com/aixcoder-plugin/aiXcoder-7B/tree/main/aiXcoder_ex-tended_dataset provided by [11] to evaluate the model’s Fill-in-the-Middle (FIM) code completion performance on files from real-world open-source projects. The specific evaluation metrics used were Exact Match (EM) and Edit Similarity (ES). Our implementation is based on the reproduction project of Qwen Coder [^30] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval. We filtered out test cases where the total length of the context and the expected completion exceeded 8192 tokens or where the generation length was greater than 64 tokens. During evaluation, the temperature was set to 0, and the maximum output length of the model was capped at 64 tokens.
@@
As shown in Table II, SynthCoder-8B achieves an improvement of over 10% after finetuning, surpassing all baseline models, including those with a larger total parameter count such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT demonstrates an overall improvement of more than 10% compared to the base model, with a maximum gain of 22% on the JavaScript EM score. The aligned SynthCoder-8B-DPO-wo-RS shows further modest gains across all metrics, improving upon SynthCoder-8B-SFT by 4.2% in overall EM and 1.6% in overall ES. The final model with repetition suppression, SynthCoder-8B-DPO-RS, trades a slight decrease of up to 1.2% in C++ ES and on Python metrics for a 2.1% increase in overall EM when compared to SynthCoder-8B-DPO-wo-RS. Ultimately, this results in a total improvement of 17.8% in overall EM and 12.3% in overall ES relative to the base model.
@@
Seed-Coder-8B-Inst 38.6 $\downarrow$28.2% 72.0 $\downarrow$12.8% 33.1 $\downarrow$37.4% 67.2 $\downarrow$17.2% 44.0 $\downarrow$28.4% 74.5 $\downarrow$15.3% 38.6 $\downarrow$16.9% 70.5 $\downarrow$11.2% 39.0 $\downarrow$27.6% 71.3 $\downarrow$14.3%
Seed-Coder-8B-Base 44.9 $\downarrow$10.2% 75.3 $\downarrow$7.8% 35.0 $\downarrow$30.0% 66.9 $\downarrow$17.8% 46.7 $\downarrow$20.9% 75.3 $\downarrow$14.1% 41.6 $\downarrow$8.4% 72.5 $\downarrow$8.3% 42.3 $\downarrow$17.8% 72.6 $\downarrow$12.3%
SynthCoder-8B-SFT 46.3 $\downarrow$6.7% 79.3 $\downarrow$2.4% 42.1 $\downarrow$8.0% 77.5 $\downarrow$1.6% 52.8 $\downarrow$6.9% 83.9 $\downarrow$2.4% 43.7 $\downarrow$3.2% 78.3 $\downarrow$0.1% 46.9 $\downarrow$6.3% 80.2 $\downarrow$1.7%
SynthCoder-8B-DPO-wo-RS 48.8 $\downarrow$1.4% 81.4 $\uparrow$0.3% 43.5 $\downarrow$4.7% 78.5 $\downarrow$0.4% 55.1 $\downarrow$2.4% 85.2 $\downarrow$0.8% 45.2 $\uparrow$0.2% 79.4 $\uparrow$1.2% 48.8 $\downarrow$2.1% 81.5 $\downarrow$0.1%
SynthCoder-8B-DPO-RS 49.5 81.2 45.5 78.8 56.5 85.9 45.1 78.5 49.8 81.5
IV-B Repository Level Code Completion
We selected ExecRepoBench, CrossCodeEval, and CoLT to evaluate the LLM’s capabilities in repository-level code completion.
IV-B1 CrossCodeEval with BM25 Retrieval Strategy
CrossCodeEval is a benchmark designed to evaluate the ability of LLMs to leverage cross-file context for code completion. Its data is sourced from open-source GitHub projects and covers four programming languages: Python, Java, TypeScript, and C#. The benchmark employs static analysis methods to construct test cases that are solvable only with the aid of cross-file context [14]. In this paper, we adopt the same evaluation metrics as the original paper – EM, ES, and ES_repoeval (abbreviated as ES_R) – and use the official cross-file context data provided by CrossCodeEval, which is retrieved using the BM25 text similarity algorithm. Each retrieved code snippet is limited to a maximum of 10 lines, and the top 5 most relevant snippets are selected. For each model, we utilize its specific prompt format for completion, with a input context length of 8192 tokens, a maximum output length of 50 tokens, and a temperature of 0.
As shown in Table III, SynthCoder-8B demonstrates a modest overall improvement compared to its base model and achieves the best performance among all models tested. This includes models with a larger parameter count, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT shows an overall increase of 1.1% in EM, 0.6% in ES, and 0.7% in ES_R. While there was a slight decrease in the metrics for Java, all other languages showed improvement. The aligned SynthCoder-8B-DPO-wo-RS achieves gains of 0.6% to 0.7% overall, and also secures the highest scores on Java within the SynthCoder-8B series. The final version with repetition suppression, SynthCoder-8B-DPO-RS, obtains slight improvements of 1.5%, 1.5%, and 1.1% in overall EM, ES, and ES_R, respectively, when compared to the base model.

### v3

@@ III-B3 Rejected Code Sampling
To ensure the quality and diversity of the negative samples, we apply a series of filters to the candidates. First, validity filtering is performed, which includes deduplication, removal of null values, and elimination of samples that are identical to or contain the ground truth. Subsequently, similarity filtering based on BLEU scores is used to remove negative samples that are too similar to the ground truth. After the screening process, we retain up to 3 negative samples for each original data point, which constitute the paired dataset for preference alignment training.
@@
To suppress the model’s tendency to simply repeat the context, we randomly sampled 10% of the data from rejected code sampling where the expected completion does not start with the first line of the suffix. We constructed suffix-repetition suppression pairs where the positive example is the ground-truth completion and the negative example is the first line of the suffix code. This data was then added to the synthesized dataset. Similarly, we also added 1% of prefix-repetition suppression data.
@@ III-C Model Post-training
During both training stages, the SPM (Suffix-Prefix-Middle) FIM prompt format is adopted, as described in its technical report [9]. In reference of [17], we set FIM rate to 0.7.
@@ III-C1 Finetune Stage
The data preprocessing for curriculum learning fine-tuning is divided into two steps:
  • Step 1: Complexity Sorting.
    We use Tree-sitter to parse code files and generate Abstract Syntax Trees (ASTs), then measure complexity by the number of identifier nodes in the AST and sort them in descending order as is described in [14]. Symbols are extracted by recursively traversing the ASTs.
  • Step 2: Sample Selection.
@@
In the curriculum learning fine-tuning stage, we performed full-parameter fine-tuning on the Seed-Coder-8B-Base model, updating all of its weights. The model was trained for one epoch using the AdamW optimizer with a learning rate of $3\times 10^{-7}$ and a 50-step linear warmup. The training data consisted of the top 30% ($k=0.3$) most complex samples from the original dataset, with all input sequences truncated to a maximum length of 8192 tokens. The training was conducted on four NVIDIA H800 (80GB PCIe) GPUs, using a global batch size of 256 and a micro-batch size of 1 per GPU. To enhance efficiency, we employed several optimization strategies, including DeepSpeed ZeRO-1 [44], Flash Attention [45], and gradient checkpointing [^9] https://github.com/cybertronai/gradient-checkpointing. This fine-tuning process was completed in 800 steps over approximately 20 hours. Checkpoints are recorded every 100 steps and only the best checkpoint will be retained.
@@ III-C2 Alignment Stage
In the preference alignment stage using the DPO algorithm, the DPO parameter $\beta$ was set to 0.9, with an equivalent batch size of 64. The model was trained for one epoch with a maximum learning rate of 1e-4, a linear learning rate warmup of 32 steps, and a cosine learning rate decay scheduler. The maximum length for each training data was 8192 tokens. For parameter-efficient fine-tuning, we employed the Low-Rank Adaptation (LoRA) technique [47]. The LoRA rankrank was set to 32 and the scaling factor alphaalpha was set to 64. The training was conducted on 4 NVIDIA H800 80G PCIE GPUs, leveraging the TRL framework [^10] https://github.com/huggingface/trl (version 0.17.0) in conjunction with DeepSpeed ZeRO-2 [44] (with a CPU offloading strategy) and gradient checkpointing. The changes in Loss, Reward Accuracy, and Reward Margin during the training process are shown in Figure 2. The alignment process lasted for 400 steps within 24 hours, and it can be observed that the Training Loss decreased relatively steadily, while the Reward Accuracy and Reward Margin increased and then remained relatively stable. Checkpoints are recorded every 100 steps and only the best checkpoint will be retained.
@@ IV Evaluation
We conducted a comprehensive and rigorous evaluation of SynthCoder and baseline models on both file-level and repository-level code completion tasks.
IV-A Evaluation Setup
IV-A1 Benchmark Baselines
Due to constraints in computational resources, we selected representative open-source models with Fill-in-the-Middle (FIM) capabilities with size ranging from 6B to 8B that were released between 2024 and the first half of 2025. Models released earlier, such as CodeLlama-7b-hf [^11] https://huggingface.co/codellama/CodeLlama-7b-hf from 2023, were excluded. The final set of models used for comparison is as follows.
  • aiXcoder-7B-V2121212https://huggingface.co/aiXcoder/aiXcoder-7B-V2 A 7B parameter code LLM designed to fully leverage long-context information, particularly excelling at cross-file code completion.
  • StarCoder2-7B [^13] https://huggingface.co/bigcode/starcoder2-7b A 7B parameter code LLM developed by the BigCode project.
  • DeepSeek-Coder-6.7B (-base [^14] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-base, -instruct [^15] https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct) A pair of 6.7B code LLMs developed by DeepSeek, while deepseek-coder-6.7b-instruct is the instruction-tuned version of deepseek-coder-6.7b-base.
  • Qwen2.5-Coder-7B [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-7B (-Instruct [^17] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct) A pair of 7B code LLMs developed by Alibaba, while Qwen2.5-Coder-7B-Instruct is the instruction-tuned version.
  • Seed-Coder-8B (-Base [^18] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Base, -Instruct [^19] https://huggingface.co/ByteDance-Seed/Seed-Coder-8B-Instruct) A pair of 8B parameter code LLMs developed by the ByteDance, while Seed-Coder-8B-Instruct is the instruction-tuned version of Seed-Coder-8B-Base.
Additionally, we include Mixture-of-Experts (MoE) models that, despite having fewer than 10B active parameters, possess a total parameter count that exceeds this threshold.
@@ IV-A1 Benchmark Baselines
    A pair of lightweight code LLMs in DeepSeek-Coder-V2 model family [^22] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Base, while DeepSeek-Coder-V2-Lite-Instruct is the instruction-tuned version.
@@
    An efficient instruction-tuned code model from Alibaba’s Qwen3 Coder series.
To ensure the rigor of our experimental results, we only included models in our comparison that were evaluated in the aforementioned file-level FIM code completion benchmarks and for which the official developers provided a specific FIM prompt format in their technical reports or open-source repository documentation. Note that the aiXcoder-7B-V2 model’s technical report [10] and open-source repository [^24] https://github.com/aixcoder-plugin/aiXcoder-7B-V2 only provide a prompt format for file-level completion and do not specify how to pass cross-file context. Therefore, we only evaluated it in the file-level code completion tasks.
IV-A2 SynthCoder Variants for Ablation Study
To validate the impact of different training stages and data compositions, we evaluated three versions of our model: SynthCoder-8B-SFT, SynthCoder-8B-DPO-wo-RS, and SynthCoder-8B-DPO (the final version, referred to as SynthCoder-8B). We compared these against the base model, Seed-Coder-8B-Base. The differences between these model versions are detailed in Table I.
@@ IV-A2 SynthCoder Variants for Ablation Study
SynthCoder-8B-DPO YES YES YES
IV-A3 Benchmark Datasets
We select aiXcoder dataset as the evaluation dataset to verify our model’s capacity on file level code completetion, and ExecRepoBench, CrossCodeEval, and CoLT to evaluate the LLM’s capabilities in repository-level code completion, as is shown below. There’s always been only one trial for each model in all the benchmarks respectively. All models were deployed on a single local H800 80G PCIe GPU, and inference was performed using VLLM version 0.9.1252525https://github.com/vllm-project/vllm. For all the generation process, we employed deterministic decoding with the following hyperparameters: $temperature=0$, top_p=1.0top\_p=1.0, frequency_penalty=0.0frequency\_penalty=0.0, repetition_penalty=1.0repetition\_penalty=1.0, and length_penalty=1.0length\_penalty=1.0.
aiXcoder
@@ IV-A3 Benchmark Datasets
The aiXcoder dataset [11] evaluates Fill-in-the-Middle (FIM) code completion on real-world open-source projects. Our evaluation follows the Qwen Coder reproduction script [^26] https://github.com/QwenLM/Qwen3-Coder/tree/main/qwencoder-eval. We filtered out test cases where the combined length of the context and ground truth exceeded 8192 tokens, or where the ground truth itself was longer than 64 tokens. During inference, we used a temperature of 0 and capped the maximum generation length at 64 tokens.
CrossCodeEval with BM25 Retrieval
CrossCodeEval [19] is a benchmark designed to assess cross-file code completion. It uses static analysis to create test cases from projects in Python, Java, TypeScript, and C# that are only solvable with external context. For our evaluation, we use the officially provided context retrieved via the BM25 algorithm. This context includes the top 5 most relevant code snippets, each truncated to a maximum of 10 lines. We adhere to each model’s specific prompt format, setting the maximum input length to 8192 tokens and the maximum output to 50 tokens.
CoLT with Dependency-based Retrieval
The CoLT benchmark [10] assesses cross-file context utilization across 12,000 samples in Python, Java, C/C++, and Go. To specifically evaluate the model’s ability to leverage code structure, we exclusively used the context retrieved through dependency call relationships. For each model, we followed its native prompt format, standardizing the maximum input length to 4096 tokens and the maximum output length to 200 tokens.
ExecRepoBench
ExecRepoBench [^27] https://huggingface.co/datasets/CSJianYang/ExecRepoBench evaluates code completion within real-world Python repositories [12]. Uniquely, it provides test cases with both local, intra-file context in a Fill-in-the-Middle (FIM) format and relevant cross-file context as complete files. In our evaluation, we implemented a rule-based post-processing step to correct indentation errors in the generated code. The intra-file and cross-file contexts were each limited to 4096 tokens, and the maximum generation length was set to 160 tokens.
IV-B Precision in Benchmarks
The precision of our model on the benchmark datasets is reported in this section. The evaluation is based on the following metrics.
  • Exact Match (EM): The model generation result is identical with the canonical solution. Adopted in aiXcoder, CrossCodeEval and CoLT.
  • Edit Similarity (ES): The edit similarity between model generation result and canonical solution calculated by the FuzzyWuzzy library  [^28] https://github.com/seatgeek/fuzzywuzzy. Adopted in aiXcoder, CrossCodeEval, CoLT and ExecRepoBench.
  • Edit Similarity RepoEval (ES_R): The edit similarity between model generation result and canonical solution calculated in the way of RepoEval [20]. Adopted only in CrossCodeEval.
  • Pass@1: Whether the complete code assembled by the prefix, suffix and generation results can pass all the corresponding test cases. Adopted only in ExecRepoEval.
To conserve space in the data tables, we abbreviate Instruct to Inst and DeepSeek to DS. To save space, when presenting the precision metrics, all numerical values in tables are rounded to one decimal place, and all the percentages are calculated between current varient of SynthCoder and the final SynthCoder-8B-DPO while blue means that the latter is better.
IV-B1 aiXcoder
@@ IV-B Precision in Benchmarks
As shown in Table II, SynthCoder-8B-SFT achieves an improvement of over 10% after finetuning, surpassing all baseline models, including those with a larger total parameter count such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT demonstrates an overall improvement of more than 10% compared to the base model, with a maximum gain of 20.3% on the JavaScript EM score. The aligned SynthCoder-8B-DPO-wo-RS shows further modest gains across all metrics, improving upon SynthCoder-8B-SFT by 4.1% in overall EM and 1.6% in overall ES. The final model with repetition suppression, SynthCoder-8B-DPO, trades a slight decrease of up to 1.2% in Python ES for a 2.1% increase in overall EM when compared to SynthCoder-8B-DPO-wo-RS. Ultimately, this results in a total improvement of 17.8% in overall EM and 12.3% in overall ES relative to the base model, showing the effectiveness of our approach.
@@
Seed-Coder-8B-Inst 38.6 $\uparrow$28.2% 72.0 $\uparrow$12.8% 33.1 $\uparrow$37.4% 67.2 $\uparrow$17.2% 44.0 $\uparrow$28.4% 74.5 $\uparrow$15.3% 38.6 $\uparrow$16.9% 70.5 $\uparrow$11.2% 39.0 $\uparrow$27.6% 71.3 $\uparrow$14.3%
Seed-Coder-8B-Base 44.9 $\uparrow$10.2% 75.3 $\uparrow$7.8% 35.0 $\uparrow$30.0% 66.9 $\uparrow$17.8% 46.7 $\uparrow$20.9% 75.3 $\uparrow$14.1% 41.6 $\uparrow$8.4% 72.5 $\uparrow$8.3% 42.3 $\uparrow$17.8% 72.6 $\uparrow$12.3%
SynthCoder-8B-SFT 46.3 $\uparrow$6.7% 79.3 $\uparrow$2.4% 42.1 $\uparrow$8.0% 77.5 $\uparrow$1.6% 52.8 $\uparrow$6.9% 83.9 $\uparrow$2.4% 43.7 $\uparrow$3.2% 78.3 $\uparrow$0.1% 46.9 $\uparrow$6.3% 80.2 $\uparrow$1.7%
SynthCoder-8B-DPO-wo-RS 48.8 $\uparrow$1.4% 81.4 $\downarrow$0.3% 43.5 $\uparrow$4.7% 78.5 $\uparrow$0.4% 55.1 $\uparrow$2.4% 85.2 $\uparrow$0.8% 45.2 $\downarrow$0.2% 79.4 $\downarrow$1.2% 48.8 $\uparrow$2.1% 81.5 $\uparrow$0.1%
SynthCoder-8B-DPO 49.5 81.2 45.5 78.8 56.5 85.9 45.1 78.5 49.8 81.5
IV-B2 CrossCodeEval with BM25 Retrieval Strategy
As shown in Table III, SynthCoder-8B-DPO demonstrates a modest overall improvement compared to its base model, and generally achieves the best performance among all models tested. This includes models with a larger parameter count, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. Specifically, the fine-tuned SynthCoder-8B-SFT shows an overall increase of 2.1% in EM, 0.9% in ES, and 0.6% in ES_R. While there was a slight decrease in the metrics for Java, all other languages showed improvement. The aligned SynthCoder-8B-DPO-wo-RS achieves gains of 0.6% to 0.7% overall, and also secures the highest scores on Java within the SynthCoderseries. The final version with repetition suppression, SynthCoder-8B-DPO, obtains slight improvements of 1.5%, 1.5%, and 1.1% in overall EM, ES, and ES_R, respectively, when compared to the base model Seed-Coder-8B-Base.